# Active Learning with Task–Driven Representations for Messy Pools

### v1

@@ 1 Introduction
Active learning (MacKay, 1992b, ; Settles,, 2009) is a framework for selecting the best data points to label during the training of a predictive model. It has the potential to be especially useful in the setting of messy, uncurated pools commonly encountered with real–world data, where the unlabeled data points have widely varying relevance to our task of interest (Bickford Smith et al.,, 2023, 2024; Sun et al.,, 2017; Citovsky et al.,, 2021). For example, our pool may contain many examples of classes that we are not interested in predicting or the subtle variations we are trying to pick up on may be dwarfed by irrelevant features.
@@
Unfortunately, such messiness can undermine standard active learning pipelines. In particular, simple predictive uncertainty measures will often be highest for irrelevant points in the pool (Bickford Smith et al.,, 2023). Previous work has looked to address this by developing acquisition functions that are more robust to forms of messiness like class imbalance and irrelevant datapoints (Zhang et al.,, 2022; Kothawade et al.,, 2021; Zhang et al.,, 2023; Nuggehalli et al.,, 2023; Xie et al.,, 2024).
@@
The success of active learning methods, though, is critically dependent not only on our acquisition strategy, but our model choice as well. In particular, there is a growing body of evidence that it is imperative to use semi–supervised models to incorporate the rich information available in the unlabeled data, for the sake of both direct prediction and guiding acquisition (Burkhardt et al.,, 2018; Hacohen et al.,, 2022; Seo et al., 2022a, ; Mittal et al.,, 2023; Sener and Savarese,, 2017; Yehuda et al.,, 2022; Lüth et al.,, 2023; Chan et al.,, 2021; Bhatnagar et al.,, 2020; Ebrahimi et al.,, 2020; Gao et al.,, 2020). By contrast, the aforementioned approaches for dealing with messy pools have all focused mainly on fully supervised models. Bickford Smith et al., (2024) recently provided a first notable exception to this by showing that their prediction-orientated acquisition strategy can be successfully combined with unsupervised representations to yield state-of-the-art active learning performance for messy pools.
@@
In this work, we show that effectively dealing with messy pools requires careful specific considerations in how our model is constructed, not just our acquisition strategy. In particular, we highlight that the current use of unsupervised representations can itself undermine our ability to effectively deal with messy pools: the task-agnostic nature of such representations mean that they can fail to capture all the information relevant to our task. As a result, the representations can fail to capture the right notion of similarity between inputs for our task, leading to inaccurate predictive correlations and, ultimately, suboptimal acquisitions.
@@
We suggest to address this issue by using task–driven representations. Namely, we argue that updating our representations at (semi–)regular intervals during the active learning process allows us to guide the representations towards capturing task–relevant information. This, in turn, enables our model to better learn the relevant similarities between inputs, improves uncertainty estimation, and ultimately leads to better acquisition decisions.
@@
To this end, we introduce two concrete strategies for learning these task–driven representations. The first is to periodically retrain the representation using a semi–supervised objective based on the original pool data and the labels collected thus far. Specifically, we build on the CCVAE approach of Joy et al., (2021) to learn representations where a subset of the latents are guided by a downstream classifier to capture information in the labels. The second is a more lightweight approach where we periodically perform a simple supervised fine–tuning of the original unsupervised representation. Empirically, we find that both approaches lead to more effective acquisitions and significantly enhance model performance.
@@ 2.1 Problem Formulation
Active learning (AL) provides a principled approach to adaptively selecting datapoints to label when training a predictive model. We consider pool–based AL where we have a small initial labeled dataset $\mathcal{D}_{l}=\{(x^{l}_{i},y^{l}_{i})\}_{i=1}^{M}$ and a larger unlabeled pool $\mathcal{D}_{u}=\{(x^{u}_{i})\}_{i=1}^{N}$, with inputs $x\in\mathcal{X}$ and outputs $y\in\mathcal{Y}$. The objective is to iteratively choose (a subset of) points from $\mathcal{D}_{u}$ for labeling, with the aim of producing the best model with the fewest labels.
@@ 2.2 Active Learning with Messy Pools
In real–world data, the unlabeled data points in our pool can have widely varying relevance to our task of interest. Pools of web–scraped audio, images and text are common examples of this. Active learning ought to be particularly helpful in dealing with this messiness, by identifying only the most useful inputs to label. However, it can cause problems for many standard active learning pipelines, as predictive uncertainty is often highest on these irrelevant datapoints (Bickford Smith et al.,, 2023, 2024).
Previous work in this setting has primarily focused on dealing with such messiness by designing appropriate acquisition functions. Notable works include SIMILAR (Kothawade et al.,, 2021) which uses submodular information measures as acquisition functions to deal with pools involving class imbalance and redundant classes; GALAXY (Zhang et al.,, 2022) which proposes a graph–based acquisition function that has shown state–of–the–art performance on pools with redundant and imbalanced classes; and DIRECT (Nuggehalli et al.,, 2023) which uses a boundary-aware, one-dimensional acquisition strategy to deal with both class imbalance and label noise.
The acquisition strategy we will primarily utilise in this work is EPIG (Bickford Smith et al.,, 2023), which uses a prediction–oriented acquisition strategy to deal with imbalanced and redundant classes. Specifically, EPIG introduces a target input distribution $p_{*}(x_{*})$ and then considers the expected uncertainty reduction (Rainforth et al.,, 2024; Lindley,, 1956) in hypothetical future predictions $y_{*}|x_{*}$:
@@
where H refers to Shannon entropy (Shannon,, 1948). By focusing on a particular target prediction task, EPIG allows acquisitions to be focused on datapoints that will aid downstream performance and hopefully avoid irrelevant points in the pool.
@@ 2.3 Semi–Supervised Active Learning
Previous work on active learning with messy pools has primarily used fully supervised models (Kothawade et al.,, 2021; Zhang et al.,, 2022; Nuggehalli et al.,, 2023; Bickford Smith et al.,, 2023; Xie et al.,, 2024). This is in spite of a growing line of work which shows that it is typically imperative to use semi–supervised models for most active learning problems (Burkhardt et al.,, 2018; Hacohen et al.,, 2022; Seo et al., 2022a, ; Mittal et al.,, 2023): by incorporating the rich information available in the unlabeled data, semi–supervised approaches can improve immediate predictive performance and also the quality of uncertainty estimates that drive acquisitions.
Recently, Bickford Smith et al., (2024) observed reliable gains over random acquisition, in messy and non-messy settings, by first learning unsupervised representations from the unlabeled data and then performing active learning on top of those fixed representations with a fully supervised prediction head. Concretely, they decompose the predictive model $p_{\phi}(y|x)$ into a fixed deterministic encoder $g:X\rightarrow\mathbb{R}^{d}$ and a stochastic prediction head $p_{\phi}(y|z,\theta_{h})$, where $z=g(x)$ and $\theta_{h}\sim p_{\phi}(\theta_{h})$. The overall predictive model is then given by $p_{\phi}(y|x)=\mathbb{E}_{p_{\phi}(\theta_{h})}\left[p_{\phi}(y|g(x),\theta_{h})\right]$. By fixing the encoder while updating only the prediction head between active learning iterations, they are able to leverage large encoders pretrained on the pool that capture much of the information needed for the downstream task in a lower-dimensional latent space (Chen et al., 2020b, ; Chen et al., 2020a, ), while using smaller prediction heads that improve the computational efficiency of the active learning and the quality of the updates.
@@ 3 Shortfalls of Unsupervised Representations for Messy Pools
We now explain why dealing with messy pools requires careful consideration of the underlying model setup and not just the choice of acquisition function. Namely, we explain how using unsupervised representations, as per (Bickford Smith et al.,, 2024), can itself break down in the messy pool scenario.
@@
To this end, we first highlight three key features that can occur with real–world pools and be problematic for active learning approaches: i) class imbalance, an uneven distribution of classes; ii) redundant classes, where the pool contains datapoints that are not one of the classes we wish to predict (e.g. our pool is images of animals but our task is specifically classifying dog breeds); and iii) redundant information, where there is significant information in the data beyond that is irrelevant to what we are trying to predict (e.g. detecting legions in MRI brain data, where variations in equipment used to make scans causes large irrelevant variations in datapoints). These “messiness” characteristics are present in a wide range of scenarios ranging from web-scraped data to natural data (Kim et al.,, 2023; Ren et al.,, 2023; Ardila et al.,, 2019; Gemmeke et al.,, 2017) and also encompass most forms of data messiness considered in prior work (Zhang et al.,, 2022; Kothawade et al.,, 2021; Xie et al.,, 2024; Zhang et al.,, 2023; Nuggehalli et al.,, 2023).
@@
A weakness of unsupervised representations here is that as our data becomes increasingly messy, the representations may fail to capture all the information relevant for our task. This has been observed outside of the active learning context for various unsupervised representation learning methods (Caron et al.,, 2019; Tian et al.,, 2021; Shi et al.,, 2022; Huang et al.,, 2022). At a high level, it comes from unsupervised representations being task–agnostic: as the pool becomes messier, the task–specific information becomes smaller compared to the task–irrelevant information, and the representation increasingly focuses on the latter. Even if the task–specific information has not been lost completely by the unsupervised representation, its dilution will generally still increase the difficulty of learning an effective prediction head  (Cole et al.,, 2022; Huang et al.,, 2022).
@@
A direct consequence of this is that it will inevitably hurt the performance of active learning algorithms. We observe this in Figure 1 for the case of EPIG, where we perform active learning on a balanced pool, but with unsupervised representations that were pre–trained on pools of increasing messiness. This is expected, as here selecting the most informative data points relies on the model’s ability to make accurate similarity judgments in the latent space (Bickford Smith et al.,, 2024). Capturing these similarities is essential for establishing the predictive correlations that drive effective exploration and exploitation in active learning (Wang et al.,, 2021; Osband et al., 2022b, ; Osband et al., 2022c, ). However, these similarities are task–dependent and break down with messier pools as our representations fail to include relevant task–specific information.
@@ 4 Using Task–Driven Representations
Motivated by the issues discussed in Section 3, we propose to instead use task-driven representations for active learning with messy pools. Our suggested approach builds on the semi–supervised approach of Bickford Smith et al., (2024) described in Section 2. However, instead of using a fixed unsupervised encoder, we regularly update it as we acquire more labels using a semi–supervised representation learning technique. That is, our predictive model is given by $p_{\phi}(y|{x})=\mathbb{E}_{p(\theta_{h})}\left[p_{\phi}(y|g(x),\theta_{h})\right],$ where $p_{\phi}(y|z,\theta_{h})$ is our prediction head with stochastic parameters $\theta_{h}\sim p_{\phi}(\theta_{h})$ and $z=g(x)$ are our representations as before, but $g:X\rightarrow\mathbb{R}^{d}$ is now a semi–supervised encoder that utilises both the unlabeled data and acquired labels.
There are a variety of different methods one could use to learn this task-driven semi–supervised encoder (e.g. Kingma et al., (2014); Narayanaswamy et al., (2017); Chen et al., 2020b ; Assran et al., (2021); Mo et al., (2023)). Something they generally have in common is that they utilise a “guidance classifier”, $c:\mathbb{R}^{d}\to[0,1]^{|\mathcal{Y}|}$, that maps representations to class probabilities. This classifier will either be learned alongside the encoder itself, typically by maximising an objective that accounts for both fidelity of the representation across all the data and the performance of the classifier on the labelled data, or will be used as part of a fine–tuning procedure to update a pretrained unsupervised encoder. The aim of this is to guide the representations to be task–driven, such that they retain the information required for both effective downstream prediction and label acquisition. To complement our task–driven representations, we use the EPIG acquisition function for our approach, which benefits from treating the guidance classifier and prediction head as distinct components, as we explain later.
The best setup to use for training/updating the encoder will inevitably vary between problems. Below, we outline two possible concrete approaches. The first is inspired by the CCVAE approach of Joy et al., (2021) and involves fully retraining the encoder using a semi–supervised variational objective that encourages the label information to concentrate in a small subset of the learned latents. This has the advantage of providing a low-dimensional representation that is strongly tailored to the task, but the retraining can be expensive. The second is a more lightweight approach that simply uses supervised fine–tuning of the original unsupervised representation. This has the advantage of speed and simplicity, but may make it harder to balance pool and label information in the representation.
@@ 4.1 A Split Representation Approach
Encoder training  Unlike in the original CCVAE, we have no need to perform generations or interventions with our representation. We therefore eschew the introduction of an additional conditional generative model on $z_{c}|y$ and directly train the encoder and downstream classifier in an end–to–end manner using both the labeled and unlabeled data. Specifically, we maximise the following objective, corresponding to Equation (2) in Joy et al., (2021),
@@
where $\mathcal{L}(\lambda,\psi;x)=\mathbb{E}_{q_{\lambda}(z|x)}\left[\log\left(p_{\psi}(x|z)p(z)/q_{\lambda}(z|x)\right)\right]$ is the standard VAE objective (Kingma et al.,, 2013), $q_{\lambda}(z|x)$ is the VAE encoder with parameters $\lambda$ (and we take $g(x)=\mathbb{E}_{q_{\lambda}(z|x)}[z]$), $p_{\psi}(x|z)$ is the VAE decoder with parameters $\psi$, $p(z)$ is a fixed isotropic Gaussian prior, $c_{\omega}$ is the downstream classifier with parameters $\omega$, $\mathcal{D}_{\mathrm{pool}}$ is the unlabelled pool data, $\mathcal{D}_{\mathrm{labelled}}$ is the labelled data gathered thusfar, and $\alpha$ is a hyperparameter controlling the label pressure on $z_{c}$.
Following Joy et al., (2021), we perform the optimization using stochastic gradient ascent where updates with the labelled and unlabelled data are conducted in separate batches. As semi–supervised encoders typically struggle with class imbalance and the low–data regimes considered in active learning (Oliver et al.,, 2018; Yu et al.,, 2020; Guo et al.,, 2020), we further perform simple data augmentations on our labelled set and upsample minority classes. To deal with the redundant classes in our pool, we follow Bickford Smith et al., (2023, 2024) by labelling them as a single “redundant” category and retaining them in our labelled set, noting that these labels still contain useful information for future acquisition by marking points as not being one of the target classes (Yang et al.,, 2023).
@@
The prediction head, on the other hand, needs to be probabilistic with well–calibrated reducible uncertainty estimates. It will be updated at every iteration so it should ideally be cheap to update, and it should not require careful hyperparamter tuning or access to validation data. In our experiments, we use Random Forests (Breiman,, 2001), due to their fast training and strong “out–of–the–box” performance, and ablate with their different prediction heads in the Appendix.
Encoder retraining   We retrain our encoder regularly after every $k$ acquired labels. We recommend using larger values of $k$ ($\gtrsim 25$ in our experiments) to keep computational costs lows and because very small choices of $k$, and in particular taking $k=1$, has the potential to harm performance, by creating a disconnect between the update strategy assumed by the acquisition function (which is based only on the prediction head) and the actual updates performed.

### v2

@@ 1 Introduction
Active learning (MacKay, 1992b; Settles, 2009) is a framework for selecting the best datapoints to label during the training of a predictive model. It has the potential to be especially useful in the setting of messy, uncurated pools commonly encountered with real–world data, where the unlabelled data points have widely varying relevance to our task of interest (Bickford Smith et al., 2023, 2024; Sun et al., 2017; Citovsky et al., 2021). For example, our pool may contain many examples of classes that we are not interested in predicting or the subtle variations we are trying to pick up on may be dwarfed by irrelevant features.
@@
Unfortunately, such messiness can undermine existing active learning pipelines. In particular, simple predictive uncertainty measures will often be highest for irrelevant points in the pool (Bickford Smith et al., 2023). Previous work has tried to address this by developing acquisition functions that are more robust to forms of messiness like class imbalance and irrelevant data points (Zhang et al., 2022; Kothawade et al., 2021; Zhang et al., 2023; Nuggehalli et al., 2023; Xie et al., 2024).
@@
The success of active learning methods, though, is critically dependent not only on our acquisition strategy, but our model choice as well. In particular, there is a growing body of evidence that it is imperative to use semi–supervised models to incorporate the rich information available in the unlabelled data, for the sake of both direct prediction and guiding acquisition (Burkhardt et al., 2018; Hacohen et al., 2022; Seo et al., 2022b; Mittal et al., 2023; Sener & Savarese, 2017; Yehuda et al., 2022; Lüth et al., 2023; Chan et al., 2021; Bhatnagar et al., 2020; Ebrahimi et al., 2020; Gao et al., 2020). By contrast, the aforementioned approaches for dealing with messy pools have all focused mainly on fully supervised models. Bickford Smith et al. (2024) recently provided a first notable exception to this by showing that their prediction-orientated acquisition strategy can be successfully combined with unsupervised representations to yield state-of-the-art active learning performance for messy pools.
@@
In this work, we show that effectively dealing with messy pools requires careful specific considerations in how our model is constructed, not just our acquisition strategy. In particular, we highlight that the current use of unsupervised representations can itself undermine our ability to effectively deal with messy pools: the task-agnostic nature of such representations mean that they can fail to capture all the information relevant to our task. As a result, the representations can fail to capture the right notion of similarity between inputs for our task, leading to inaccurate predictive correlations and, ultimately, suboptimal acquisitions (Wang et al., 2021; Osband et al., 2022b).
@@
We suggest to address this issue by using task–driven representations. Namely, we argue that updating our representations at regular intervals during the active learning process allows us to guide the representations towards capturing task–relevant information. This, in turn, enables our model to better learn the relevant similarities between inputs, improves uncertainty estimation, and ultimately leads to better acquisition decisions.
@@
To this end, we introduce two concrete strategies for learning these task–driven representations. The first is to periodically retrain the representation using a semi–supervised objective based on the original pool data and the labels collected thus far. Specifically, we build on the CCVAE approach of Joy et al. (2021) to learn representations where a subset of the latents are guided by a downstream classifier to capture information in the labels. The second is a more lightweight approach where we periodically perform a simple supervised fine–tuning of the original unsupervised representation. Empirically, we find that both approaches lead to more effective acquisitions and significantly enhance model performance.
@@ 2.1 Problem Formulation
Active learning (AL) provides a principled approach to adaptively selecting datapoints to label when training a predictive model. We consider pool–based AL where we have a small initial labelled dataset $\mathcal{D}_{l}=\{(x^{l}_{i},y^{l}_{i})\}_{i=1}^{M}$ and a larger unlabelled pool $\mathcal{D}_{u}=\{(x^{u}_{i})\}_{i=1}^{N}$, with inputs $x\in\mathcal{X}$ and outputs $y\in\mathcal{Y}$. The objective is to iteratively choose (a subset of) points from $\mathcal{D}_{u}$ for labeling, with the aim of producing the best model with the fewest labels.
@@ 2.2 Active Learning with Messy Pools
In real–world data, the unlabelled data points in our pool can have widely varying relevance to our task of interest. Pools of web–scraped audio, images and text are common examples of this. Active learning ought to be particularly helpful in dealing with this messiness, by identifying only the most useful inputs to label. However, it can cause problems for many standard active learning pipelines, as predictive uncertainty is often highest on these irrelevant data points (Bickford Smith et al., 2023, 2024).
Previous work in this setting has primarily focused on dealing with such messiness by designing appropriate acquisition functions. Notable works include SIMILAR (Kothawade et al., 2021) which uses submodular information measures as acquisition functions to deal with pools involving class imbalance and redundant classes; GALAXY (Zhang et al., 2022) which proposes a graph–based acquisition function that has shown state–of–the–art performance on pools with redundant and imbalanced classes; and DIRECT (Nuggehalli et al., 2023) which uses a boundary-aware, one-dimensional acquisition strategy to deal with both class imbalance and label noise.
The acquisition strategy we will primarily utilise in this work is EPIG (Bickford Smith et al., 2023), which uses a prediction–oriented acquisition strategy to deal with imbalanced and redundant classes. Specifically, EPIG introduces a target input distribution $p_{*}(x_{*})$ and then considers the expected uncertainty reduction (Rainforth et al., 2024; Lindley, 1956) in hypothetical future predictions $y_{*}|x_{*}$:
@@
where H refers to Shannon entropy (Shannon, 1948). By focusing on a particular target prediction task, EPIG allows acquisitions to be focused on data points that will aid downstream performance and hopefully avoid irrelevant points in the pool.
@@ 2.3 Semi–Supervised Active Learning
Previous work on active learning with messy pools has primarily used fully supervised models (Kothawade et al., 2021; Zhang et al., 2022; Nuggehalli et al., 2023; Bickford Smith et al., 2023; Xie et al., 2024). This is in spite of a growing line of work which shows that it is typically imperative to use semi–supervised models for most active learning problems (Burkhardt et al., 2018; Hacohen et al., 2022; Seo et al., 2022b; Mittal et al., 2023): by incorporating the rich information available in the unlabelled data, semi–supervised approaches can improve immediate predictive performance and also the quality of uncertainty estimates that drive acquisitions.
Recently, (Bickford Smith et al., 2024) observed reliable gains over random acquisition, in messy and non-messy settings, by first learning unsupervised representations from the unlabelled data and then performing active learning on top of those fixed representations with a fully supervised prediction head. Concretely, they decompose the predictive model $p_{\phi}(y|x)$ into a fixed deterministic encoder $g:X\rightarrow\mathbb{R}^{d}$ and a stochastic prediction head $p_{\phi}(y|z,\theta_{h})$, where $z=g(x)$ and $\theta_{h}\sim p_{\phi}(\theta_{h})$. The overall predictive model is then given by $p_{\phi}(y|x)=\mathbb{E}_{p_{\phi}(\theta_{h})}\left[p_{\phi}(y|g(x),\theta_{h})\right]$. By fixing the encoder while updating only the prediction head between active learning iterations, they are able to leverage large encoders pretrained on the pool that capture much of the information needed for the downstream task in a lower-dimensional latent space (Chen et al., 2020b, a), while using smaller prediction heads that improve the computational efficiency of the active learning and the quality of the updates.
@@ 3 Shortfalls of Unsupervised Representations for Messy Pools
We now explain why dealing with messy pools requires careful consideration of the underlying model setup, not just the choice of acquisition function. Namely, we explain how using unsupervised representations, as per (Bickford Smith et al., 2024), can itself break down with messy pools.
@@
To this end, we first highlight three key features that can occur with real–world pools and be problematic for active learning approaches: i) class imbalance, an uneven distribution of classes; ii) redundant classes, where the pool contains data points that are not one of the classes we wish to predict (e.g. our pool is images of animals but our task is specifically classifying dog breeds); and iii) redundant information, where there is significant information in the data that is irrelevant to what we are trying to predict (e.g. detecting legions in MRI brain data, where variations in equipment used to make scans causes large irrelevant variations in data points). These “messiness” characteristics are present in a wide range of scenarios ranging from web-scraped data to natural data (Kim et al., 2023; Ren et al., 2023; Ardila et al., 2019; Gemmeke et al., 2017) and also encompass most forms of data messiness considered in prior work (Zhang et al., 2022; Kothawade et al., 2021; Xie et al., 2024; Zhang et al., 2023; Nuggehalli et al., 2023).
@@
A weakness of unsupervised representations here is that as our data becomes increasingly messy, the representations may fail to capture all the information relevant for our task. This has been observed outside of the active learning context for various unsupervised representation learning methods (Caron et al., 2019; Tian et al., 2021; Shi et al., 2022; Huang et al., 2022). At a high level, it comes from unsupervised representations being task–agnostic: as the pool becomes messier, the task–specific information becomes smaller compared to the task–irrelevant information, and the representation increasingly focuses on the latter. Even if the task–specific information has not been lost completely by the unsupervised representation, its dilution will generally still increase the difficulty of learning an effective prediction head  (Cole et al., 2022; Huang et al., 2022).
@@
A direct consequence of this is that it will inevitably hurt the performance of active learning algorithms. We observe this in Figure 1 for the case of EPIG, where we perform active learning on a balanced pool, but with unsupervised representations that were pre–trained on pools of increasing messiness. This is expected, as here selecting the most informative data points relies on the model’s ability to make accurate similarity judgments in the latent space (Bickford Smith et al., 2024). Capturing these similarities is essential for establishing the predictive correlations that drive effective exploration and exploitation in active learning (Wang et al., 2021; Osband et al., 2022b, c). However, these similarities are task–dependent and break down with messier pools as our representations fail to include relevant task–specific information.
@@ 4 Using Task–Driven Representations
Motivated by the issues discussed in §3, we propose to instead use task-driven representations for active learning with messy pools. Our suggested approach builds on the semi–supervised approach of Bickford Smith et al. (2024) described in §2. However, instead of using a fixed unsupervised encoder, we regularly update it as we acquire more labels using a semi–supervised representation learning technique. That is, our predictive model is given by $p_{\phi}(y|{x})=\mathbb{E}_{p(\theta_{h})}\left[p_{\phi}(y|g(x),\theta_{h})\right],$ where $p_{\phi}(y|z,\theta_{h})$ is our prediction head with stochastic parameters $\theta_{h}\sim p_{\phi}(\theta_{h})$ and $z=g(x)$ are our representations as before, but $g:X\rightarrow\mathbb{R}^{d}$ is now a semi–supervised encoder that utilises both the unlabelled data and acquired labels.
There are a variety of different methods one could use to learn this task-driven semi–supervised encoder (e.g. (Kingma et al., 2014; Narayanaswamy et al., 2017; Chen et al., 2020b; Assran et al., 2021; Mo et al., 2023)). Something they generally have in common is that they utilise a “guidance classifier”, $c:\mathbb{R}^{d}\to[0,1]^{|\mathcal{Y}|}$, that maps representations to class probabilities. This classifier will either be learned alongside the encoder itself, typically by maximising an objective that accounts for both fidelity of the representation across all the data and the performance of the classifier on the labelled data, or will be used as part of a fine–tuning procedure to update a pretrained unsupervised encoder. The aim of this is to guide the representations to be task–driven, such that they retain the information required for both effective downstream prediction and label acquisition. To complement our task–driven representations, we use the EPIG acquisition function for our approach, which benefits from treating the guidance classifier and prediction head as distinct components, as we explain later.
The best setup to use for training/updating the encoder will inevitably vary between problems. Below, we outline two possible concrete approaches. The first is inspired by the CCVAE approach of Joy et al. (2021) and involves fully retraining the encoder using a semi–supervised variational objective that encourages the label information to concentrate in a small subset of the learned latents. This has the advantage of providing a low-dimensional representation that is strongly tailored to the task, but the retraining can be expensive. The second is a more lightweight approach that simply uses supervised fine–tuning of the original unsupervised representation. This has the advantage of speed and simplicity, but may make it harder to balance pool and label information in the representation.
@@ 4.1 A Split Representation Approach
Encoder training  Unlike in the original CCVAE, we have no need to perform generations or interventions with our representation. We therefore eschew the introduction of an additional conditional generative model on $z_{c}|y$ and directly train the encoder and downstream classifier in an end–to–end manner using both the labelled and unlabelled data. Specifically, we maximise the following objective, corresponding to Equation (2) in Joy et al. (2021),
@@
where $\mathcal{L}(\lambda,\psi;x)=\mathbb{E}_{q_{\lambda}(z|x)}\left[\log\left(p_{\psi}(x|z)p(z)/q_{\lambda}(z|x)\right)\right]$ is the standard VAE objective (Kingma et al., 2013), $q_{\lambda}(z|x)$ is the VAE encoder with parameters $\lambda$ (and we take $g(x)=\mathbb{E}_{q_{\lambda}(z|x)}[z]$), $p_{\psi}(x|z)$ is the VAE decoder with parameters $\psi$, $p(z)$ is a fixed isotropic Gaussian prior, $c_{\omega}$ is the downstream classifier with parameters $\omega$, $\mathcal{D}_{\mathrm{pool}}$ is the unlabelled pool data, $\mathcal{D}_{\mathrm{labelled}}$ is the labelled data gathered thusfar, and $\alpha$ is a hyperparameter controlling the label pressure on $z_{c}$.
Following Joy et al. (2021), we perform the optimization using stochastic gradient ascent where updates with the labelled and unlabelled data are conducted in separate batches. As semi–supervised encoders typically struggle with class imbalance and the low–data regimes considered in active learning (Oliver et al., 2018; Yu et al., 2020; Guo et al., 2020), we further perform simple data augmentations on our labelled set and upsample minority classes. To deal with the redundant classes in our pool, we follow Bickford Smith et al. (2023, 2024) by labelling them as a single “redundant” category and retaining them in our labelled set, noting that these labels still contain useful information for future acquisition by marking points as not being one of the target classes (Yang et al., 2023).
@@
The prediction head, on the other hand, needs to be probabilistic with well–calibrated reducible uncertainty estimates. It will be updated at every iteration so it should ideally be cheap to update, and it should not require careful hyperparamter tuning or access to validation data. In our experiments, we use Random Forests (Breiman, 2001), due to their fast training and strong “out–of–the–box” performance, and ablate with their different prediction heads in the Appendix.
Encoder retraining   We retrain our encoder regularly after every $k$ acquired labels. We recommend using larger values of $k$ ($\gtrsim 25$ in our experiments) to keep computational costs low and because very small choices of $k$, and in particular taking $k=1$, has the potential to harm performance, by creating a disconnect between the update strategy assumed by the acquisition function (which is based only on the prediction head) and the actual updates performed.